# Dynamics of Two Species with Density-Dependent Interactions in a Mutualistic Context

### v1

@@ 3 General dynamics of two species with any types of interactions
In this work, we study the behavior of system (1), including stability, oscillations, and other dynamics, with a particular focus on how the specific forms of $f$ and $g$ influence species coexistence. In Section 3.1 we consider a general framework in which $f$ and $g$ are assumed to be continuous only, and we apply index theory (see Perko, (2013) and Appendix A for details) to prove Theorem 3.1. Then, using our general Theorem 3.1, we study the different assumptions described in Section 2.2 adding the assumption that $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$ to investigate more precisely how the nature of the interactions considered can affect the outcomes of the system. We mainly focus on simple equilibrium points and use methods based on the Jacobian matrix to analyze their stability and nature. For example, regardless of the specific nature of the interaction represented by system (1), certain information can already be deduced about the equilibrium point $\left(0,0\right)$ by applying these methods (see Appendix B and Brauer and Castillo-Chavez, (2012)).
3.1 Assumptions and Equilibrium Index Alternation
We introduce several assumptions about the functions $f$ and $g$ that are commonly derived from biological models of population dynamics. Most of these assumptions, inspired by models reviewed in Hale and Valdovinos, (2021), are robust and remain valid even when $f$ and $g$ are subject to minor local perturbations. We recall that the zeros of $f$ and $g$ define the isoclines of the system, and their intersections determine the equilibrium points.
@@ 3.1 Assumptions and Equilibrium Index Alternation
These assumptions directly influence the properties of the vector field. One characteristic of equilibrium points is their index, which describes the local behavior of the flow in their neighborhood. The index helps determine whether an equilibrium point acts as an attractive, a repulsive, or a saddle point, and reveals how trajectories interact with the point (see Perko, (2013) and Appendix A for more details). In particular, for systems satisfying H1 to H4 on $f$ and $g$, the indices of equilibrium points in the positive quadrant exhibit a regular alternation along the isoclines. This result is formalized in the following theorem:
@@
We recall that the model cover the wide parasitic–mutualistic continuum. The interactions may therefore represent competition, predation, or mutualism, as long as the functions $f$ and $g$ satisfy the conditions stated in assumptions H1H4.
@@ 3.2 Intraspecific Competition and Cooperation Assumption
In many models presented in Hale and Valdovinos, (2021), intraspecific competition is a key mechanism in regulating population dynamics within each species, preventing the uncontrolled population growth which may be criticized in some models of mutualistic interactions. With this additional assumption, all previous equilibrium points with an index of +1+1 become attractive points.
@@
For a model satisfying the conditions of Corollary 3.1, if there are at least two equilibrium points in the positive quadrant, then an attractive point exists. This implies the existence of conditions under which coexistence of the two species is possible.
Under the opposite constant sign assumptions, i.e. intraspecific cooperation for both species as described in 2.2, the system in the positive quadrant alternates between repulsive nodes and saddle points along the isoclines.
@@
Thus, negative assumptions on the sign of the partial derivatives $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$ (which biologically correspond to intraspecific competition, always negative regardless of species densities) prevent the emergence of any coexistence solutions other than attractive equilibrium points or saddle points.
3.3 Illustrative Examples of Theorem 3.1
In this section, we illustrate Theorem 3.1 using four models of mutualistic interactions reviewed in Hale and Valdovinos, (2021), each represented on Figure 4 (in the order presented), each emphasizing a different characteristic of mutualism. These four models use different mathematical representations to emphasize certain aspects of mutualism over others, and describe facultative (Neuhauser and Fargione,, 2004) or obligate (Graves et al.,, 2006), density-dependent, or saturating mutualistic interactions. Although their mathematical formulation reflects different perspectives on mutualistic interactions, as discussed in Section 2.2, we show that they satisfy the assumptions of system (1), highlighting the generality of the framework provided in our study.
@@
    In Zhang, (2003) the author models the dynamics of two mutualistic species using the same mathematical formulation twice, reflecting an interaction between species at the same trophic level. In this model, the growth rate of each species is influenced by a quadratic term that depends on the density of the other species. The intraspecific competition is represented by a constant term, while the interaction between species varies in sign: it remains positive up to a threshold, beyond which it becomes negative. This formulation captures the effect of saturation and illustrates the continuum between mutualism and parasitism. The system is given by:
    $\begin{cases}\dot{x}=r_{x}x\left(c_{x}-x-a_{x}(y-b_{x})^{2}\right),\\ \dot{y}=r_{y}y\left(c_{y}-y-a_{y}(x-b_{y})^{2}\right),\end{cases}$ (3)
    where $r_{x}$, $r_{y}$, $a_{x}$, $a_{y}$, $c_{x}$, $c_{y}$ are positive real constants and $b_{x}$, $b_{y}$ are real constants. Here, there is the assumption of intraspecific competition, where $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$. Moreover, the interaction is mutualistic as long as $y$ remains below the threshold $b_{x}$, but when $y$ exceeds this value, $\frac{\partial f}{\partial y}$ turns negative. This dynamic highlights how the strength and nature of the interaction shift depending on species densities.
@@
    Neuhauser and Fargione, (2004) introduce a model describing facultative mutualism between plants ($x$) and mycorrhizal fungi ($y$). In this formulation, the effects on species $y$ remain constant, regardless $x$ or $y$, with intraspecific competition always negative ($\frac{\partial g}{\partial y}<0$) and the mutualistic benefit from $x$ always positive ($\frac{\partial g}{\partial x}>0$). Similarly, intraspecific competition for $x$ is constant and negative ($\frac{\partial f}{\partial x}<0$). However, the term $\frac{\partial f}{\partial y}$, which characterizes the nature of the interaction, varies depending on the densities of both species. Specifically, as $y$ increases, the benefit for $x$ diminishes and can eventually turn negative, shifting the interaction away from mutualism. However, this effect is modulated by the presence of $x$: for huge values of $x$, $\frac{\partial f}{\partial y}$ is more strongly positive, reinforcing mutualistic benefits. $x$ and $y$ must increase together to prevent the relationship from becoming parasitic. The system is given by:
    $\begin{cases}\dot{x}=r_{x}x\left(\frac{K_{x}\gamma_{xy}y-x}{K_{x}+\gamma_{xy}y}-ay\right),\\ \dot{y}=r_{y}y\left(\frac{K_{y}+\alpha_{yx}x-y}{K_{y}}\right),\end{cases}$ (4)
    where all parameters are positive real constants.
@@
    The model proposed by Graves et al., (2006) represents obligate mutualism for lichens, where the benefits of the interaction are tightly linked to the densities of both species. The system assumes constant negative intraspecific competition, with identical functional forms for both species, and the mutualistic interaction varies with the density of the other species. Specifically, as the density of $y$ increases, the effect on $x$ decreases, but without any saturation effect—just a continuous increase with density.
    The system is given by:
    $\begin{cases}\dot{x}=x\left(r_{x0}+(r_{x1}-r_{x0})\left(1-\exp{(-k_{x}y)}\right)-a_{x}x\right),\\ \dot{y}=y\left(r_{y0}+(r_{y1}-r_{y0})\left(1-\exp{(-k_{y}x)}\right)-a_{y}y\right).\end{cases}$ (5)
    where $r_{x1}$, $r_{y1}$, $a_{x}$, $a_{y}$, $k_{x}$, $k_{y}$ are positive real constants and $r_{x0}\leq r_{x1}$, $r_{y0}\leq r_{y1}$, $r_{x0}$ and $r_{y0}$ are real constants. Here, the interaction benefits are regulated by an exponential term, which prevents unbounded growth as species densities increase.
@@
    A more general consumer-resource approach is presented by Holland and DeAngelis, (2010), which models the bidirectional exchange of benefits in plant-mycorrhizal systems. In this framework, one species provides resources while the other enhances nutrient uptake efficiency. The system follows a similar functional form for both species, but the dynamics are more complex as all aspects—including the intraspecific competition—vary with the densities of both species.
    The system is given by:
    $\begin{cases}\dot{x}=x\left(r_{x}+c_{x}\left(\frac{a_{xy}y}{h_{y}+y}\right)-q_{x}\left(\frac{\beta_{xy}y}{e_{x}+x}\right)-s_{x}x\right),\\ \dot{y}=y\left(r_{y}+c_{y}\left(\frac{a_{yx}x}{h_{x}+x}\right)-q_{y}\left(\frac{\beta_{yx}x}{e_{y}+y}\right)-s_{y}y\right).\end{cases}$ (6)
    where all parameters are positive constants, except $r_{x}$ and $r_{y}$, which are real constants. In this model, the mutualistic interaction is obligate when the intrinsic growth rates are negative ($r_{x}\leq 0$, $r_{y}\leq 0$). The interaction structure is dynamic, with the intraspecific competition for $x$ decreasing as $y$ grows, but increasing as $x$ increases. Conversely, the effect of $y$ on $x$ increases with $x$, but decreases as $y$ increases. This creates a more flexible framework for describing how species’ interactions evolve with changing densities.
@@
Figure 4: Phase portraits of four differents models of mutualism. Each model highlights different aspects of mutualism while satisfying the assumptions H1 to H4 of system (1) and exhibiting an alternation of equilibrium point indices along the isoclines. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. White points represent repulsive points, black points represent attractive points and gray points represent saddle points. (4(a)) Parameters for Zhang, (2003) model, in (3): $c_{x}=7$, $c_{y}=8$, $a_{x}=0.7$, $a_{y}=0.5$, $b_{x}=4.5$, $b_{y}=3$, $r_{x}=7$, $r_{y}=4$. (4(b)) For Neuhauser and Fargione, (2004) model, in (4): $r_{x}=1.4$, $r_{y}=1$, $K_{x}=2$, $K_{y}=5$, $\gamma_{xy}=4$, $\alpha_{xy}=1.5$, $a=0.2$. (4(c)) For Graves et al., (2006) model, in (5): $r_{x0}=-2$, $r_{y0}=0.5$, $r_{x1}=4$, $r_{y1}=5$, $k_{x}=0.2$, $k_{y}=0.2$, $a_{x}=0.35$, $a_{y}=0.5$. (4(d)) For Holland and DeAngelis, (2010) model, in (6) : $r_{x}=2$, $r_{y}=3$, $c_{x}=3$, $c_{y}=1$, $a_{xy}=1$, $a_{yx}=2.6$, $h_{x}=1$, $h_{y}=1$, $q_{x}=1$, $q_{y}=2.7$, $\beta_{xy}=1$, $\beta_{yx}=1.5$, $e_{x}=0.5$, $e_{y}=3$, $s_{x}=0.7$, $s_{y}=0.5$.
@@ 4 Periodic solutions in the dynamics of two mutualistic species
A system of the form (1) can represent either predator-prey, competition, or mutualistic interactions. The results described so far do not distinguish between these different types of interactions.
@@
Limit cycles differ from centers, such as those observed in the Lotka-Volterra predator-prey system. Centers are equilibrium points where the trajectories of the system form closed, periodic orbits, but they are neither attractive nor repulsive. As a result, the behavior of the system is highly sensitive to initial conditions: each new set of initial conditions produces a distinct trajectory, making the system highly sensitive to small variations (May,, 1972). However, even in the case of centers, when the Lotka-Volterra predator-prey model is modified to represent mutualistic interactions, the centers disappear.
@@ 4.2 Case of extended mutualism
In this section, we use the extended definition of mutualism provided in Definition 2.1, which includes the previous case of strict mutualism and describes partially mutualistic interactions, where the interactions between species can be mutualistic only for certain combinations of species densities. As previously discussed, the emergence of a limit cycle can indeed only occur in a context where the interaction is either parasitic or competitive, and not in a region where both $\frac{\partial f}{\partial x}>0$ and $\frac{\partial g}{\partial y}>0$ simultaneously and remain unchanged. Here, we consider density-dependent effects that give rise to regions where the mutualistic interaction locally shifts to a parasitic one. We identify two different scenarios where the nature of the interaction is not fixed, i.e., where $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ change signs, leading to the emergence of a limit cycle.
@@
    Figure 5: Phase portrait assuming extended mutualism leading to a cyclic behaviour (illustration of 1.10 of the rectangular region $ABCD$). The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. By assumption 1.9, $(x^{*},y^{*})$ is a repulsive point (the white point), which implies, by Theorem 3.1, that $(x_{2},0)$ is a saddle point (the grey points), as well as $(0,0)$.
Remark 4.3.
These assumptions are more restrictive than those in Section 3.1 and imply assumptions H1 to H4.
Under conditions 1.1-1.10, we have the following result, which will be proven in Section 6.
Theorem 4.3.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying conditions 1.1-1.10. Then a limit cycle exists inside the positive quadrant.
Example 4.1.
Let us consider the following system:
Figure 6: Phase portrait leading to a limit cycle. Simulation of system (7) with parameters: $a_{1}=3.7$, $b_{1}=0.5$, $c_{1}=2$, $d_{1}=0.6$ and $a_{2}=3$, $b_{2}=0.2$, $c_{2}=4$, $d_{2}=1$. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. The white point represents a repulsive point and gray points represent saddle points. The purple trajectory starts at a population density of 2 for species $x$ and 6 for species $y$. It converges toward the limit cycle from the outside, moving away from the extinction states, either of both species or the extinction of $y$. The gray trajectory starts at densities of 7 for species $x$ and 10 for species $y$, and converges toward the limit cycle from the inside, moving away from the interior equilibrium point.
@@
As before, these assumptions are more restrictive than those presented in Section 3.1 and imply assumptions H1 to H4.
Under conditions 2.1-2.10, we have the following result, which will be proven in Section 6.
Theorem 4.4.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying conditions 2.1-2.10. Then a cycle limit exists inside the positive quadrant.

### v2

@@ 3 General dynamics of two species with any types of interactions
In this work, we study the behavior of system (1), including stability, oscillations, and other dynamics, with a particular focus on how the specific forms of $f$ and $g$ influence species coexistence. In Sections 3.1, 3.2 and 3.3 we consider a general framework in which $f$ and $g$ are assumed to be continuous only, and we apply index theory (see Perko (2013) and Appendix A for details) to prove Theorem 3.1. Then, in Section 3.4, using our general Theorem 3.1, we more precisely investigate how the nature of the interactions can affect the outcomes of the system under the additional assumption that $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$. We mainly focus on simple equilibrium points and use methods based on the Jacobian matrix to analyze their stability and nature. For example, regardless of the specific nature of the interaction represented by system (1), certain information can already be deduced about the equilibrium point $\left(0,0\right)$ by applying these methods (see Appendix B and Brauer and Castillo-Chavez (2012)).
3.1 Assumptions
We introduce several assumptions about the functions $f$ and $g$ that are commonly derived from biological models of population dynamics. Most of these assumptions, inspired by models reviewed in Hale and Valdovinos (2021), are robust and remain valid even when $f$ and $g$ are subject to minor local perturbations. We recall that the zeros of $f$ and $g$ define the isoclines of the system, and their intersections determine the equilibrium points.
@@ 3.1 Assumptions
These assumptions directly influence the properties of the vector field.
3.2 Equilibrium Index Alternation
One characteristic of equilibrium points is their index, which describes the local behavior of the flow in their neighborhood. The index helps determine whether an equilibrium point acts as an attractive, a repulsive, or a saddle point, and reveals how trajectories interact with the point (see Perko (2013) and Appendix A for more details). In particular, for systems satisfying H1 to H4 on $f$ and $g$, the indices of equilibrium points in the positive quadrant exhibit a regular alternation along the isoclines. This result is formalized in the following theorem:
@@ 3.2 Equilibrium Index Alternation
We recall that the model covers the wide parasitic–mutualistic continuum. The interactions may therefore represent competition, predation, or mutualism, as long as the functions $f$ and $g$ satisfy the conditions stated in assumptions H1H4.
@@
3.3 Generalisation to Axes
Theorem 3.1 describes the alternation of indices for equilibrium points located strictly inside the positive quadrant, i.e. at intersections of $\Gamma_{f}$ and $\Gamma_{g}$ with strictly positive coordinates. However, equilibrium points may also lie on the axes. These additional equilibrium points may arise from the intersections of $\Gamma_{f}$ with the axis $x=0$, or of $\Gamma_{g}$ with the axis $y=0$. For population models, such equilibria correspond to situations in which one species persists at equilibrium while the other is extinct.
For isoclines $\Gamma_{f}$ and $\Gamma_{g}$ satisfying hypotheses H1 to H4, a finite number of geometric configurations may occur along the coordinate axes. Since each isocline is continuous and consists of a single connected component, it can intersect the axes at most twice. Taking into account the relative ordering of these intersections along the axes, a total of $71$ distinct configurations are possible. Among them, $18$ configurations generate at least one equilibrium point located on an axis and are therefore not covered by Theorem 3.1. Under the same geometric assumptions on the isoclines, this alternation property can be extended to include such boundary equilibria.
Proposition 3.5.
Under the assumptions of Theorem 3.1, the alternation of equilibrium point indices along the isoclines extends to include equilibrium points located on the coordinate axes, with the exception of the origin $(0,0)$.
This result shows that the index alternation is a global geometric property of the vector field induced by (1), and is not restricted to strictly positive coexistence equilibria. Equilibrium points located on the axes naturally fit into the same alternation pattern, while the origin $(0,0)$ plays a singular role and is therefore excluded from this statement.
@@ 3.4 Intraspecific Competition and Cooperation Assumption
In many models presented in Hale and Valdovinos (2021), intraspecific competition is a key mechanism in regulating population dynamics within each species, preventing the uncontrolled population growth which may be criticized in some models of mutualistic interactions. With this additional assumption, all previous equilibrium points with an index of +1+1 become attractive points.
@@
For a model satisfying the conditions of Corollary 3.6, if there are at least two equilibrium points in the positive quadrant, then an attractive point exists. This implies the existence of conditions under which the coexistence of the two species is possible.
Thus, negative assumptions on the sign of the partial derivatives $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$ (which biologically correspond to intraspecific competition, always negative regardless of species densities) prevent the emergence of any coexistence solutions other than attractive equilibrium points or saddle points. As in Proposition 3.5, this corollary can be extended to the coordinate axes, where the same monotonicity assumptions guarantee uniqueness and attractivity of equilibria.
Corollary 3.8.
Under the assumptions of Corollary 3.6, if an equilibrium point exists on a coordinate axis in $\mathbb{R}_{+}\times\mathbb{R}_{+}$, then it is unique on that axis. Moreover, if this equilibrium has index +1+1, then it is an attractive node.
Under the opposite constant sign assumptions, i.e. intraspecific cooperation for both species as described in Section 2.2, the system in the positive quadrant alternates between repulsive nodes and saddle points along the isoclines.
@@
This corollary can also be extended to the coordinate axes: the same monotonicity properties imply that any equilibrium lying on an axis is unique there, and if it has index (+1), it is a repulsive equilibrium.
Corollary 3.10.
Under the assumptions of Corollary 3.9, if an equilibrium point exists on a coordinate axis in $\mathbb{R}_{+}\times\mathbb{R}_{+}$, then it is unique on that axis. Moreover, if this equilibrium has index +1+1, then it is a repulsive node.
3.5 Illustration of Theorem 3.1 with Mutualistic Models from the Literature
In this section, we illustrate Theorem 3.1 using four models of mutualistic interactions reviewed in Hale and Valdovinos (2021) (see Table 2 and Table 3), each represented on Figure 4 (in the order presented), each emphasizing a different characteristic of mutualism. Although their mathematical formulation reflects different perspectives on mutualistic interactions, as discussed in Section 2.2, we show that they satisfy the assumptions of system (1), highlighting the generality of the framework provided in our study.
@@
    In Zhang (2003) the author models the dynamics of two mutualistic species using the same mathematical formulation twice, reflecting an interaction between species at the same trophic level. In this model, the growth rate of each species is influenced by a quadratic term that depends on the density of the other species. The intraspecific competition is represented by a constant term, while the interaction between species varies in sign: it remains positive up to a threshold, beyond which it becomes negative. This formulation captures the effect of saturation and illustrates the continuum between mutualism and parasitism. Here, there is the assumption of intraspecific competition, where $\frac{\partial f}{\partial x}<0$ and $\frac{\partial g}{\partial y}<0$. Moreover, the interaction is mutualistic as long as $y$ remains below the threshold $b_{1}$, but when $y$ exceeds this value, $\frac{\partial f}{\partial y}$ turns negative. This dynamic highlights how the strength and nature of the interaction shift depending on species densities.
@@
    Neuhauser and Fargione (2004) introduce a model describing facultative mutualism between plants ($x$) and mycorrhizal fungi ($y$). In this formulation, the effects on species $y$ remain constant, regardless $x$ or $y$, with intraspecific competition always negative $\left(\frac{\partial g}{\partial y}<0\right)$ and the mutualistic benefit from $x$ always positive $\left(\frac{\partial g}{\partial x}>0\right)$. Similarly, intraspecific competition for $x$ is constant and negative $\left(\frac{\partial f}{\partial x}<0\right)$. However, the term $\frac{\partial f}{\partial y}$, which characterizes the nature of the interaction, varies depending on the densities of both species. Specifically, as $y$ increases, the benefit for $x$ diminishes and can eventually turn negative, shifting the interaction away from mutualism. However, this effect is modulated by the presence of $x$: for huge values of $x$, $\frac{\partial f}{\partial y}$ is more strongly positive, reinforcing mutualistic benefits. $x$ and $y$ must increase together to prevent the relationship from becoming parasitic.
@@
    The model proposed by Graves et al. (2006) represents obligate mutualism for lichens, where the benefits of the interaction are tightly linked to the densities of both species. The system assumes constant negative intraspecific competition, with identical functional forms for both species, and the mutualistic interaction varies with the density of the other species. Specifically, as the density of $y$ increases, the effect on $x$ decreases, but without any saturation effect—just a continuous increase with density. Here, the interaction benefits are regulated by an exponential term, which prevents unbounded growth as species densities increase.
@@
    A more general consumer-resource approach is presented by Holland and DeAngelis (2009, 2010), which models the bidirectional exchange of benefits in plant-mycorrhizal systems. In this framework, one species provides resources while the other enhances nutrient uptake efficiency. The system follows a similar functional form for both species, but the dynamics are more complex as all aspects—including the intraspecific competition—vary with the densities of both species. In this model, the mutualistic interaction is obligate when the intrinsic growth rates are negative ($r_{1}\leq 0$, $r_{2}\leq 0$). The interaction structure is dynamic, with the intraspecific competition for $x$ decreasing as $y$ grows, but increasing as $x$ increases. Conversely, the effect of $y$ on $x$ increases with $x$, but decreases as $y$ increases. This creates a more flexible framework for describing how species’ interactions evolve with changing densities.
@@
Figure 4: Phase portraits of four differents models of mutualism. Each model highlights different aspects of mutualism while satisfying the assumptions H1 to H4 of system (1) and exhibiting an alternation of equilibrium point indices along the isoclines. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. White points represent repulsive points, black points represent attractive points and gray points represent saddle points. (4(a)) Parameters for Zhang (2003) model: $c_{1}=7$, $c_{2}=8$, $a_{1}=0.7$, $a_{2}=0.5$, $b_{1}=4.5$, $b_{2}=3$, $r_{1}=7$, $r_{2}=4$. (4(b)) For Neuhauser and Fargione (2004) model: $r_{1}=1.4$, $r_{2}=1$, $K_{1}=2$, $K_{2}=5$, $\gamma_{12}=4$, $\alpha_{12}=1.5$, $a=0.2$. (4(c)) For Graves et al. (2006) model: $r_{10}=-2$, $r_{20}=0.5$, $r_{11}=4$, $r_{21}=5$, $k_{1}=0.2$, $k_{2}=0.2$, $a_{1}=0.35$, $a_{2}=0.5$. (4(d)) For Holland and DeAngelis (2010) model: $r_{1}=2$, $r_{2}=3$, $c_{1}=3$, $c_{2}=1$, $a_{12}=1$, $a_{21}=2.6$, $h_{1}=1$, $h_{2}=1$, $q_{1}=1$, $q_{2}=2.7$, $\beta_{12}=1$, $\beta_{21}=1.5$, $e_{1}=0.5$, $e_{2}=3$, $s_{1}=0.7$, $s_{2}=0.5$.
@@ 4 Periodic solutions in the dynamics of two mutualistic species
Limit cycles differ from centers, such as those observed in the Lotka-Volterra predator-prey system. Centers are equilibrium points where the trajectories of the system form closed, periodic orbits, but they are neither attractive nor repulsive. As a result, the behavior of the system is highly sensitive to initial conditions: each new set of initial conditions produces a distinct trajectory, making the system highly sensitive to small variations (May, 1972). However, even in the case of centers, when the Lotka-Volterra predator-prey model is modified to represent mutualistic interactions, the centers disappear.
@@ 4.2 Case of extended mutualism
In this section, we use the extended definition of mutualism provided in Definition 2.1, which includes the previous case of strict mutualism and describes partially mutualistic interactions, where the interactions between species can be mutualistic only for certain combinations of species densities. As previously discussed, the emergence of a limit cycle can indeed only occur in a context where the interaction is either parasitic or competitive, and not in a region where both $\frac{\partial f}{\partial x}>0$ and $\frac{\partial g}{\partial y}>0$ simultaneously and remain unchanged. Here, we consider density-dependent effects that give rise to regions where the mutualistic interaction locally shifts to a parasitic one. We identify two different scenarios where the nature of the interaction is not fixed, i.e., where $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ change signs, leading to the emergence of a limit cycle. In a related but more specific modeling framework, Song et al. (2025) described another example of limit cycle dynamics, based on a detailed mutualism model in which variations in interaction costs and benefits are explicitly analyzed, and shown to induce or suppress cyclic behavior.
@@
    Figure 5: Phase portrait assuming extended mutualism leading to a cyclic behaviour (illustration of 1.10 of the rectangular region $ABCD$). The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. By assumption 1.9, $(x^{*},y^{*})$ is a repulsive point (the white point), which implies, by Theorem 3.1, that $(x_{2},0)$ is a saddle point (the grey point), as well as $(0,0)$.
@@
These assumptions 1.1-1.10 and 2.1-2.9 are more restrictive than those in Section 3.1 and imply assumptions H1 to H4.
Under conditions 1.1-1.10 or 2.1-2.9, we have the following result, which will be proven in Section 6.
Proposition 4.6.
Let a dynamical system in $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1), with functions $f$ and $g$ satisfying conditions 1.1-1.10 or 2.1-2.9. Then a limit cycle exists inside the positive quadrant.
Example 4.7.
Let us consider the following system:
Figure 7: Phase portrait leading to a limit cycle. Simulation of system (3) with parameters: $a_{1}=13.69$, $b_{1}=0.25$, $c_{1}=4$, $d_{1}=0.36$ and $a_{2}=9$, $b_{2}=0.04$, $c_{2}=16$, $d_{2}=1$. The curve $\Gamma_{g}$ is shown in blue, and $\Gamma_{f}$ in orange. The white point represents a repulsive point and gray points represent saddle points. The purple trajectory starts at a population density of 2 for species $x$ and 6 for species $y$. It converges toward the limit cycle from the outside, moving away from the extinction states, either of both species or the extinction of $y$. The gray trajectory starts at densities of 7 for species $x$ and 10 for species $y$, and converges toward the limit cycle from the inside, moving away from the interior equilibrium point.
In both models presented above, Assumption 1.2 (in the first configuration) and Assumption 2.2 (in the second configuration) impose a change of sign in the intraspecific interaction term. More precisely, for at least one species, the partial derivative of its per capita growth rate with respect to its own density satisfies in the first configuration $\frac{\partial f}{\partial x}>0$ at low density and $\frac{\partial f}{\partial x}<0$ at high density, and analogously for $g$ in the second configuration.
Biologically, this corresponds to the presence of an Allee effect: the species experiences positive density dependence at low population size, whereas intraspecific competition dominates at high density. This sign change induces a non-monotonic self-regulation mechanism. Our analysis suggests that such a mechanism is structurally linked to the existence of a limit cycle. In particular, in the absence of any sign change in the intraspecific terms for both species, oscillations cannot occur.
Proposition 4.8.
Let a dynamical system on $\mathbb{R}_{+}\times\mathbb{R}_{+}$ be described by (1). Assume that the functions $\frac{\partial f}{\partial x}$ and $\frac{\partial g}{\partial y}$ do not change sign in the positive quadrant. Then the system does not admit any limit cycle inside the positive quadrant.
In particular, a necessary condition for the existence of a limit cycle is that at least one species exhibits a sign change in its intraspecific effect. Biologically, this may correspond to an Allee-type positive density dependence at low density. The result, proved in Section 6, follows from a variant of the Dulac–Bendixson criterion (see Perko (2013)).